OpenStudy (anonymous):

Where is the vertex of the parabola? y = x^2 + 2x + 3 A. Above the x-axis B. On the y-axis C. On the x-axis D. Below the x-axis

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

\[ y = x^2 + 2x + 3 \] vertex is at \(x=-\frac{b}{2a}=-\frac{2}{2}=-1\) for the first coordinate

OpenStudy (anonymous):

the second coordinate is \((-1)^2+2(-1)+3=2\) making the vertex \((1,2)\) which is evidently above the \(x\) axis

OpenStudy (anonymous):

so yeah, A